Ticket Number: HEA-NOC/19980915-01 Ticket Status: OPEN
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> TCD
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 19980915 09:21:08 UTC Problem Start: 19980915 06:20 UTC+1
Ticket Closed: Problem End: 19980915 10:06 UTC+1
Routes to TCD were not being advertised.
All traffic to TCD
OSPF was restarted at TCD
Time to Fix:
3 Hours and 46 minutes Additional Notes:
Routes to TCD via MCI will be dampened for approx 30 minutes.